A FAKE SMOOTH CP^{2}#RP^{4}

DANIEL RUBERMAN RONALD J. STERN

Abstract. We show that the manifold ∗RP^{4}# ∗ CP^{2}, which is homotopy equivalent but not homeomorphic to RP^{4}#CP^{2}, is in fact smoothable.